OpenStudy (swissgirl):

is [7,10) a subset of {7,8,9,10}

OpenStudy (anonymous):

yes because 7 and 10 are present in main set also..

OpenStudy (swissgirl):

well for some odd reason its false

OpenStudy (swissgirl):

and i cant figure out y its false

OpenStudy (anonymous):

Are you sure about the brackets?

OpenStudy (anonymous):

")" in {7,10)

OpenStudy (swissgirl):

yaaa

OpenStudy (anonymous):

There is the trick. ')'

OpenStudy (swissgirl):

well it doesnt include 10 but then it shld still be a subset

OpenStudy (zarkon):

[7,10) is an interval of numbers...it includes numbers like 8.4533533453

OpenStudy (anonymous):

7 is including here and 10 is not .. It is about the brackets..

OpenStudy (anonymous):

Since [7, 10) is an interval of real numbers, it containers infinately many real numbers as @Zarkon says, so it can't be a subset of the set you are suggesting.

OpenStudy (swissgirl):

AWEESSSOOMMMEEEEEE

OpenStudy (swissgirl):

so wait

OpenStudy (swissgirl):

can u explain the difference btwn (5,8) and [5,8]?

OpenStudy (swissgirl):

Like my book doesnt explain the difference btwn the diff kinds of brackets

OpenStudy (zarkon):

\[(5,9)=\{x\in \mathbb{R}|5<x<9\}\] \[[5,9]=\{x\in \mathbb{R}|5\le x\le9\}\]

OpenStudy (swissgirl):

ok what abt (5,9] and [5,9)

OpenStudy (anonymous):

Well, (5,8) is an interval which doesn't include 5 nor 8, but [5,8] is the same interval which includes 5 and 8.